What is the time constant for charging an inductance L through a resistance R?

(a) LR

(b) L/R

(c) R/L

(d) RLC
Answer #1

For RL circuit

t=0

Applying KVL

V-(V_R + V_L) = 0

V_R = I * R

and V_L = L \frac{\mathrm{d} i}{\mathrm{d} t}

V = I*R+L * \frac{\mathrm{d} i}{\mathrm{d} t}

\frac{\mathrm{d} i}{\mathrm{d} t} =\frac{( V- i * R)}{L}

\frac{di}{V-iR}=\frac{dt}{L}

Now integrating both side

\int_{0}^{i} \frac{di}{V-iR}=\int_{0}^{t}\frac{dt}{L}

\ln (V-iR)_{0}^{i}\textrm{}=\frac{-Rt}{L}

\ln (V-iR) - \ln V =\frac{-Rt}{L}

\frac{-Rt}{L} = \ln(\frac{V-i R}{V})

taking antilog

e^{-\frac{Rt}{L}}=\frac{V-iR}{V}

V e^{-\frac{Rt}{L}}=V-iR

iR=V-V e^{-\frac{Rt}{L}}

i=\frac{V}{R} (1- e^{-\frac{Rt}{L}})

or i=\frac{V}{R} (1- e^{-\frac{t}{ \tau}})

\tau = L / R

answer is B